# references.bib

### v2

@@ 1 Introduction
With the rapid development of large language models and audio processing technology, large audio language models (LALMs) have demonstrated their superiority over conventional approaches in various speech and audio processing tasks. GPT-4o is first introduced and is pioneering the development of end-to-end speech interaction without intermediate textual conversions. Subsequently, many open-sourced LALMs [chen2025minmo, chu2024qwen2, defossez2024moshi, ding2025kimi, fang2024llama, huang2025stepaudioaqaa, huang2025step, nguyen2024spiritlminterleavedspoken, xie2024mini, xie2024mini2, xu2025qwen25omnitechnicalreport, zeng2024glm] are emerged, advancing multi-modal large language model capabilities in various speech and audio domains. Among these approaches, Qwen-Audio [chu2023qwen] and Qwen2-Audio [chu2024qwen2] perform audio analysis and generate textual responses to speech instructions. Qwen2.5-Omni [xu2025qwen25omnitechnicalreport] implements a thinker-talker architecture to enable full-duplex I/O during speech conversations. More recently, Kimi-Audio [ding2025kimi] has achieved impressive results on multiple speech and audio understanding benchmarks. In parallel, we have introduced Step-Audio [huang2025step] and Step-Audio-AQAA [huang2025stepaudioaqaa], the first LALMs to unify speech understanding and generation through discrete audio tokens at a scale of 130 billion parameters.
However, existing LALMs still face challenges in achieving natural and intelligent speech interaction. Previous LALMs such as Spirit LM [nguyen2024spiritlminterleavedspoken] and GLM-4-Voice [zeng2024glm] mainly focus on aligning the semantic information in speech inputs to text modal, neglecting the para-linguistic information which is also crucial for intentional understanding. Although LALMs including Qwen-Audio [chu2023qwen], Qwen2-Audio [chu2024qwen2] and Audio Flamingo series [ghosh2025audioflamingo2, goel2025audioflamingo3, kong2024audioflamingonovelaudio] are capable of comprehending such information, they typically generate only textual outputs and fail to further utilize this capability to produce coherent and expressive responses in speech conversations. Moreover, due to the complexities of multi-modal modeling, existing LALMs frequently suffer from hallucination and offer limited choices of timbres and speaking styles [defossez2024moshi, ding2025kimi], lacking access to real-world textual and acoustic knowledge.
@@
To address these issues and step into the next generation of multi-modal large language models, we present Step-Audio 2, an end-to-end large audio language model with industry-strength audio perception and speech interaction. Step-Audio 2 directly processes raw audio as input and outputs discrete text and audio tokens and has fewer parameters than Step-Audio [huang2025step]. Beyond capturing semantic information in speech, the model also comprehends para-linguistic and non-vocal information in audio. By leveraging chain-of-thought (CoT) reasoning and reinforcement learning (RL), Step-Audio 2 further utilizes such multi-modal information to generate expressive speech responses coherent to different conversation scenarios. To ground the model with real-world knowledge, Step-Audio 2 incorporates retrieval-augmented generation (RAG) and the capability to utilize various external tools, including web search and audio search, to provide more reliable and expressive responses. Specifically, we present an audio search as a tool unique to LALMs, enabling seamless speech retrieval via voice instructions and allowing the model to switch timbres and speaking styles based on the retrieved speech.
@@ 2.1 Speech and audio understanding
Recent advances in large language models (LLMs) [bai2023qwentechnicalreport, grattafiori2024llama3herdmodels, openai2023gpt4, openai2022chatgpt] have extended their application to a wide range of speech and audio understanding tasks, such as audio captioning, sound event detection, automatic speech recognition, audio classification, and audio-driven creative generation. A prevalent approach [chu2024qwen2, deshmukh2023pengi, gong2023joint, gong2023listen, ParalinGPT, tang2023salmonn] involves pairing speech encoders with lightweight, trainable adapters that project audio features into a textual embedding space compatible with LLMs. Building on this foundation, recent studies have further explored how to incorporate paralinguistic information such as emotion, intonation and speaker style, enabling LLMs to move beyond pure linguistic comprehension. For instance, ParalinGPT [ParalinGPT] focuses on enhancing a powerful text-based language model by integrating continuous speech embeddings, enabling it to capture paralinguistic signals such as emotion and prosody. SALMONN [tang2023salmonn] adopts a multi-modal strategy by freezing speech encoders Whisper [radford2023robust] and BEATs [BEATs], and connecting their outputs to an LLM via a window-level Q-Former, enabling joint modeling of linguistic and acoustic features. Seed-ASR [Seedasr] integrates LUISE-based speech representations with instructions and context, using context-aware SFT to capture semantic information. AudioPaLM [rubenstein2023audiopalm] combines PaLM-2 [anil2023palm2technicalreport] and AudioLM [borsos2023audiolm], unifying linguistic knowledge with paralinguistic features like speaker identity and intonation. LLM-based approaches [chu2023qwen, chu2024qwen2] increasingly rely on pretrained audio encoders such as Wav2Vec [wav2vec2], HuBERT [HuBERT], Whisper [radford2023robust], and WavLM [WavLM] to extract rich semantic representations from speech. At the same time, the extensive text knowledge and contextual reasoning capabilities stored in LLMs can provide valuable semantic guidance for understanding tasks.
@@ 2.2 Text-to-speech synthesis
Text-to-Speech (TTS) technology has made remarkable strides in recent years, evolving from traditional concatenative and statistical parametric approaches [deepvoice3, fastspeech2, tacotron2, wang2017tacotron] to codec-based TTS systems. Codec language models leverage a speech codec to extract discrete representations of speech [encodc, defossez2024moshi, ji2024wavtokenizer, snac, xin2024bigcodec, soundstream, zhang2025distinctive, zhang2023speechtokenizer] and utilize either autoregressive [ding2025kimi, zhang2023speechgpt] or masked language models [wang2024maskgct] to predict the corresponding speech tokens. These tokens are then synthesized into waveforms using codec vocoders. VALL-E [wang2023neural] marked a significant breakthrough in this area. It uses an autoregressive model to generate coarse codec codes, followed by a non-autoregressive model for the fine codes. Unlike VALL-E, which predicts acoustic tokens from phonemes and requires transcripts, SPEAR-TTS [SPEARTTS] uses a two-stage architecture with self-supervised audio prompts to clone unseen voices from just 3 seconds of speech. SparkTTS [wang2025spark] introduces BiCodec, a single-stream speech codec that encodes linguistic content as compact semantic tokens and speaker characteristics as fixed-length global tokens. Instead of relying on non-autoregressive models to predict residual discrete codes, methods like TorToiseTTS [TorToiseTTS], CosyVoice [du2024CosyVoice], CosyVoice 2 [du2024cosyvoice2], MiniMax-Speech [zhang2025minimaxspeech] and SEED-TTS [seedtts] adopt diffusion or flow-matching techniques as a second stage to reconstruct mel-spectrograms or continuous representations enriched with fine-grained acoustic and semantic details. Recent work, Kimi-Audio [ding2025kimi] combines Whisper features and semantic tokens for efficient modeling, with dual heads and a flow-matching detokenizer plus BigVGAN [lee2023bigvgan] for low-latency, expressive synthesis.
@@ 2.3 Speech-to-speech translation
Speech-to-speech translation (S2ST) is a crucial technology for eliminating communication barriers across languages. Traditional S2ST systems [wahlster2013verbmobil, wu2016google] typically adopt a cascaded pipeline consisting of automatic speech recognition (ASR), machine translation (MT), and TTS modules. Earlier studies [jia2019direct, jia2022translatotron, le2024transvip, lee2021textless] have shifted toward direct approaches that bypass intermediate textual representations, aiming for lower latency and better preservation of prosody and speaker characteristics. Two main types of direct S2ST methods have emerged, which are known as speech-to-spectrogram translation and speech-to-unit translation. Both directly generate target speech representations from the source speech without relying on textual transcriptions. A representative of the former is Translatotron [jia2019direct], the first end-to-end model to translate source speech directly into target spectrograms. Translatotron 2 [jia2022translatotron], further improves translation quality through a two-pass decoding mechanism [lee2021textless]. In contrast, speech-to-unit models predict discrete acoustic tokens rather than spectrograms, which are typically extracted using self-supervised speech encoders such as HuBERT [HuBERT] or WavLM [WavLM]. For instance, TransVIP [le2024transvip] employs a joint encoder-decoder architecture that first generates target text and residual vector quantization (RVQ) codes in the initial layer, followed by a non-causal language model that refines RVQ predictions in subsequent layers.
@@ 2.4 Speech-to-text and speech-to-speech conversation
Based on whether the LLM can directly understand and generate speech representations, existing systems can be categorized into end-to-end large audio language models and cascaded large audio language models. The former directly models audio inputs and outputs within a unified framework, while the latter relies on a modular pipeline involving separate ASR, LLM, and TTS components. Traditional speech-to-text and speech-to-speech systems typically adopt a cascaded architecture, as exemplified by AudioGPT [huang2024audiogpt] and Spoken-LLM [lin2024advancing]. However, the ASR + LLM + TTS pipeline incurs high latency and modular mismatches. This has spurred interest in unified end-to-end architectures for faster and more seamless integration. A major milestone in this direction is GPT-4o [hurst2024gpt], which supports direct end-to-end speech interaction without requiring intermediate textual conversions. Recently, several new end-to-end LALMs [defossez2024moshi, fang2024llama, xie2024mini, xie2024mini2, xu2025qwen25omnitechnicalreport] for speech-to-speech conversation have emerged. For instance, Moshi [defossez2024moshi] improves efficiency with an RQ-Transformer that generates text and audio tokens simultaneously. Similarly, Mini-Omni [xie2024mini] generates speech and text responses in parallel, following a strategy similar to MusicGen [musicgen], which enables lower first-token latency compared to interleaved generation designs. LUCY [gao2025lucylinguisticunderstandingcontrol] builds on the Mini-Omni architecture with enhancements for emotional expressiveness, naturalness, and informativeness in speech generation. It utilizes curated synthetic data and optimizes the training and decoding pipelines to handle multi-turn dialogue and function-call scenarios. Mini-Omni2 [xie2024mini] further extends Mini-Omni framework by integrating multimodal understanding and full-duplex interaction capabilities. LLaMA-Omni [fang2024llama] introduces a streaming, non-autoregressive speech decoder based on Connectionist Temporal Classification, enabling direct and efficient generation of discrete audio tokens without relying on step-by-step prediction. Freeze-Omni [wang2024freeze], on the other hand, freezes the LLM parameters during training, preserving its original capabilities while achieving low-latency speech-to-speech interaction through streaming and decoder integration. Qwen2.5-Omni [xu2025qwen25omnitechnicalreport] supports multimodal input and simultaneous text-speech output via a thinker-talker architecture, using TMRoPE for improved audio-visual synchronization through explicit temporal encoding.
@@ 3.1 Architecture
Different from our previous Step-Audio [huang2025step], Step-Audio 2 further integrates the generation of audio tokens into language modeling, achieves end-to-end audio perception and generation. As shown in Figure 3, Step-Audio 2 consists of an audio encoder, an audio adaptor, an LLM decoder and an audio detokenizer.
@@
The LLM decoder directly takes the latent audio features from the audio adaptor as input, and outputs an interleaved sequence of discrete text and audio tokens. We employ the tokenizer from CosyVoice 2 [du2024cosyvoice2] as the audio tokenizer. And the text and audio tokens are interleaved [huang2025stepaudioaqaa, huang2025step, zeng2024glm] at a fixed ratio and padded at the end to meet the ratio. The audio tokens are then extracted from the interleaved sequence and consumed by the audio detokenizer to generate the output waveform. The input audio features and output interleaved sequences are then pre-filled as the history information for the next round of conversation.
@@
Similar to Step-Audio [huang2025step] and Step-Audio-AQAA [huang2025stepaudioaqaa] , Step-Audio 2’s audio detokenizer also consists of a Flow Matching module and a HiFi-GAN [kong2020hifigangenerativeadversarialnetworks] vocoder. The Flow-Matching module generates Mel spectrograms from the output audio tokens, while the vocoder further converts the Mel spectrograms into waveforms. For Flow-Matching, we incorporate a CNN-based encoder layer after each self-attention module within the transformer block and train the model on 200,000 hours of high-quality speech. This enhancement significantly improves its Mel spectrogram reconstruction capability, leading to substantial gains in both pronunciation accuracy and timbre similarity.
Step-Audio 2 employs the same deployment infrastructure used in Step-Audio [huang2025step] and Step-Audio-AQAA [huang2025stepaudioaqaa], which includes a voice activity detection (VAD) module to filter out input speeches and achieves real-time voice conversation.
@@ 3.3 Supervised fine-tuning
We subsequently perform a large-scale, multi-task supervised fine-tuning (SFT) procedure [wei2021finetuned] to instruct the model to follow human intention in fluid conversations and master core tasks. We select audio data from open-source and proprietary data to ensure broad coverage and high quality. The model is trained on 4B tokens of text and audio data for a single epoch. And the learning rate decays from $10^{-5}$ to $10^{-6}$.
Specifically, we leverage extensive corpora such as GigaSpeech [gigaspeech], WenetSpeech [zhang2022wenetspeech10000hoursmultidomain], and other in-house data to enhance the model’s performance in multilingual and multi-dialect ASR scenarios. We reformat existing datasets for audio event classification and audio captioning, such as AudioSet [7952261] and AudioCaps [kim2019audiocaps], into speech question-answer pairs for audio understanding. To capture paralinguistic information beyond just semantics, we introduce a detailed speech captioning task and build an in-house dataset, requiring the model to generate comprehensive textual descriptions encompassing 11 paralinguistic and environmental aspects.
We employ high-quality, professionally labeled data collected in-house for TTS. We utilize the Chinese-to-English and English-to-Chinese subsets from the CoVoST 2 [wang2020covost2massivelymultilingual] dataset for speech-to-speech translation.

### v3

# StepFun Audio Team

@@ 1 Introduction
With the rapid development of large language models and audio processing technology, large audio language models (LALMs) have demonstrated their superiority over conventional approaches in various speech and audio processing tasks. GPT-4o is first introduced and is pioneering the development of end-to-end speech interaction without intermediate textual conversions. Subsequently, many open-sourced LALMs [9, 13, 16, 18, 21, 31, 32, 49, 71, 72, 74, 77] are emerged, advancing multi-modal large language model capabilities in various speech and audio domains. Among these approaches, Qwen-Audio [12] and Qwen2-Audio [13] perform audio analysis and generate textual responses to speech instructions. Qwen2.5-Omni [74] implements a thinker-talker architecture to enable full-duplex I/O during speech conversations. More recently, Kimi-Audio [18] has achieved impressive results on multiple speech and audio understanding benchmarks. In parallel, we have introduced Step-Audio [32] and Step-Audio-AQAA [31], the first LALMs to unify speech understanding and generation through discrete audio tokens at a scale of 130 billion parameters.
However, existing LALMs still face challenges in achieving natural and intelligent speech interaction. Previous LALMs such as Spirit LM [49] and GLM-4-Voice [77] mainly focus on aligning the semantic information in speech inputs to text modal, neglecting the para-linguistic information which is also crucial for intentional understanding. Although LALMs including Qwen-Audio [12], Qwen2-Audio [13] and Audio Flamingo series [24, 25, 43] are capable of comprehending such information, they typically generate only textual outputs and fail to further utilize this capability to produce coherent and expressive responses in speech conversations. Moreover, due to the complexities of multi-modal modeling, existing LALMs frequently suffer from hallucination and offer limited choices of timbres and speaking styles [16, 18], lacking access to real-world textual and acoustic knowledge.
@@
To address these issues and step into the next generation of multi-modal large language models, we present Step-Audio 2, an end-to-end large audio language model with industry-strength audio perception and speech interaction. Step-Audio 2 directly processes raw audio as input and outputs discrete text and audio tokens and has fewer parameters than Step-Audio [32]. Beyond capturing semantic information in speech, the model also comprehends para-linguistic and non-vocal information in audio. By leveraging chain-of-thought (CoT) reasoning and reinforcement learning (RL), Step-Audio 2 further utilizes such multi-modal information to generate expressive speech responses coherent to different conversation scenarios. To ground the model with real-world knowledge, Step-Audio 2 incorporates retrieval-augmented generation (RAG) and the capability to utilize various external tools, including web search and audio search, to provide more reliable and expressive responses. Specifically, we present an audio search as a tool unique to LALMs, enabling seamless speech retrieval via voice instructions and allowing the model to switch timbres and speaking styles based on the retrieved speech.
@@ 2.1 Speech and audio understanding
Recent advances in large language models (LLMs) [4, 29, 50, 51] have extended their application to a wide range of speech and audio understanding tasks, such as audio captioning, sound event detection, automatic speech recognition, audio classification, and audio-driven creative generation. A prevalent approach [13, 17, 27, 28, 48, 61] involves pairing speech encoders with lightweight, trainable adapters that project audio features into a textual embedding space compatible with LLMs. Building on this foundation, recent studies have further explored how to incorporate paralinguistic information such as emotion, intonation and speaker style, enabling LLMs to move beyond pure linguistic comprehension. For instance, ParalinGPT [48] focuses on enhancing a powerful text-based language model by integrating continuous speech embeddings, enabling it to capture paralinguistic signals such as emotion and prosody. SALMONN [61] adopts a multi-modal strategy by freezing speech encoders Whisper [53] and BEATs [10], and connecting their outputs to an LLM via a window-level Q-Former, enabling joint modeling of linguistic and acoustic features. Seed-ASR [5] integrates LUISE-based speech representations with instructions and context, using context-aware SFT to capture semantic information. AudioPaLM [57] combines PaLM-2 [2] and AudioLM [7], unifying linguistic knowledge with paralinguistic features like speaker identity and intonation. LLM-based approaches [12, 13] increasingly rely on pretrained audio encoders such as Wav2Vec [3], HuBERT [30], Whisper [53], and WavLM [11] to extract rich semantic representations from speech. At the same time, the extensive text knowledge and contextual reasoning capabilities stored in LLMs can provide valuable semantic guidance for understanding tasks.
@@ 2.2 Text-to-speech synthesis
Text-to-Speech (TTS) technology has made remarkable strides in recent years, evolving from traditional concatenative and statistical parametric approaches [52, 55, 59, 68] to codec-based TTS systems. Codec language models leverage a speech codec to extract discrete representations of speech [15, 16, 36, 60, 73, 76, 81, 82] and utilize either autoregressive [18, 80] or masked language models [67] to predict the corresponding speech tokens. These tokens are then synthesized into waveforms using codec vocoders. VALL-E [64] marked a significant breakthrough in this area. It uses an autoregressive model to generate coarse codec codes, followed by a non-autoregressive model for the fine codes. Unlike VALL-E, which predicts acoustic tokens from phonemes and requires transcripts, SPEAR-TTS [40] uses a two-stage architecture with self-supervised audio prompts to clone unseen voices from just 3 seconds of speech. SparkTTS [65] introduces BiCodec, a single-stream speech codec that encodes linguistic content as compact semantic tokens and speaker characteristics as fixed-length global tokens. Instead of relying on non-autoregressive models to predict residual discrete codes, methods like TorToiseTTS [6], CosyVoice [20], CosyVoice 2 [19], MiniMax-Speech [79] and SEED-TTS [1] adopt diffusion or flow-matching techniques as a second stage to reconstruct mel-spectrograms or continuous representations enriched with fine-grained acoustic and semantic details. Recent work, Kimi-Audio [18] combines Whisper features and semantic tokens for efficient modeling, with dual heads and a flow-matching detokenizer plus BigVGAN [46] for low-latency, expressive synthesis.
@@ 2.3 Speech-to-speech translation
Speech-to-speech translation (S2ST) is a crucial technology for eliminating communication barriers across languages. Traditional S2ST systems [62, 70] typically adopt a cascaded pipeline consisting of automatic speech recognition (ASR), machine translation (MT), and TTS modules. Earlier studies [38, 39, 44, 45] have shifted toward direct approaches that bypass intermediate textual representations, aiming for lower latency and better preservation of prosody and speaker characteristics. Two main types of direct S2ST methods have emerged, which are known as speech-to-spectrogram translation and speech-to-unit translation. Both directly generate target speech representations from the source speech without relying on textual transcriptions. A representative of the former is Translatotron [38], the first end-to-end model to translate source speech directly into target spectrograms. Translatotron 2 [39], further improves translation quality through a two-pass decoding mechanism [45]. In contrast, speech-to-unit models predict discrete acoustic tokens rather than spectrograms, which are typically extracted using self-supervised speech encoders such as HuBERT [30] or WavLM [11]. For instance, TransVIP [44] employs a joint encoder-decoder architecture that first generates target text and residual vector quantization (RVQ) codes in the initial layer, followed by a non-causal language model that refines RVQ predictions in subsequent layers.
@@ 2.4 Speech-to-text and speech-to-speech conversation
Based on whether the LLM can directly understand and generate speech representations, existing systems can be categorized into end-to-end large audio language models and cascaded large audio language models. The former directly models audio inputs and outputs within a unified framework, while the latter relies on a modular pipeline involving separate ASR, LLM, and TTS components. Traditional speech-to-text and speech-to-speech systems typically adopt a cascaded architecture, as exemplified by AudioGPT [33] and Spoken-LLM [47]. However, the ASR + LLM + TTS pipeline incurs high latency and modular mismatches. This has spurred interest in unified end-to-end architectures for faster and more seamless integration. A major milestone in this direction is GPT-4o [34], which supports direct end-to-end speech interaction without requiring intermediate textual conversions. Recently, several new end-to-end LALMs [16, 21, 71, 72, 74] for speech-to-speech conversation have emerged. For instance, Moshi [16] improves efficiency with an RQ-Transformer that generates text and audio tokens simultaneously. Similarly, Mini-Omni [71] generates speech and text responses in parallel, following a strategy similar to MusicGen [14], which enables lower first-token latency compared to interleaved generation designs. LUCY [22] builds on the Mini-Omni architecture with enhancements for emotional expressiveness, naturalness, and informativeness in speech generation. It utilizes curated synthetic data and optimizes the training and decoding pipelines to handle multi-turn dialogue and function-call scenarios. Mini-Omni2 [71] further extends Mini-Omni framework by integrating multimodal understanding and full-duplex interaction capabilities. LLaMA-Omni [21] introduces a streaming, non-autoregressive speech decoder based on Connectionist Temporal Classification, enabling direct and efficient generation of discrete audio tokens without relying on step-by-step prediction. Freeze-Omni [66], on the other hand, freezes the LLM parameters during training, preserving its original capabilities while achieving low-latency speech-to-speech interaction through streaming and decoder integration. Qwen2.5-Omni [74] supports multimodal input and simultaneous text-speech output via a thinker-talker architecture, using TMRoPE for improved audio-visual synchronization through explicit temporal encoding.
@@ 3.1 Architecture
Different from our previous Step-Audio [32], Step-Audio 2 further integrates the generation of audio tokens into language modeling, achieves end-to-end audio perception and generation. As shown in Figure 3, Step-Audio 2 consists of an audio encoder, an audio adaptor, an LLM decoder and an audio detokenizer.
@@
The LLM decoder directly takes the latent audio features from the audio adaptor as input, and outputs an interleaved sequence of discrete text and audio tokens. We employ the tokenizer from CosyVoice 2 [19] as the audio tokenizer. And the text and audio tokens are interleaved [31, 32, 77] at a fixed ratio and padded at the end to meet the ratio. The audio tokens are then extracted from the interleaved sequence and consumed by the audio detokenizer to generate the output waveform. The input audio features and output interleaved sequences are then pre-filled as the history information for the next round of conversation.
@@
Similar to Step-Audio [32] and Step-Audio-AQAA [31] , Step-Audio 2’s audio detokenizer also consists of a Flow Matching module and a HiFi-GAN [42] vocoder. The Flow-Matching module generates Mel spectrograms from the output audio tokens, while the vocoder further converts the Mel spectrograms into waveforms. For Flow-Matching, we incorporate a CNN-based encoder layer after each self-attention module within the transformer block and train the model on 200,000 hours of high-quality speech. This enhancement significantly improves its Mel spectrogram reconstruction capability, leading to substantial gains in both pronunciation accuracy and timbre similarity.
Step-Audio 2 employs the same deployment infrastructure used in Step-Audio [32] and Step-Audio-AQAA [31], which includes a voice activity detection (VAD) module to filter out input speeches and achieves real-time voice conversation.
@@ 3.3 Supervised fine-tuning
We subsequently perform a large-scale, multi-task supervised fine-tuning (SFT) procedure [69] to instruct the model to follow human intention in fluid conversations and master core tasks. We select audio data from open-source and proprietary data to ensure broad coverage and high quality. The model is trained on 4B tokens of text and audio data for a single epoch. And the learning rate decays from $10^{-5}$ to $10^{-6}$.
Specifically, we leverage extensive corpora such as GigaSpeech [8], WenetSpeech [78], and other in-house data to enhance the model’s performance in multilingual and multi-dialect ASR scenarios. We reformat existing datasets for audio event classification and audio captioning, such as AudioSet [23] and AudioCaps [41], into speech question-answer pairs for audio understanding. To capture paralinguistic information beyond just semantics, we introduce a detailed speech captioning task and build an in-house dataset, requiring the model to generate comprehensive textual descriptions encompassing 11 paralinguistic and environmental aspects.
We employ high-quality, professionally labeled data collected in-house for TTS. We utilize the Chinese-to-English and English-to-Chinese subsets from the CoVoST 2 [63] dataset for speech-to-speech translation.